A boomerang is thrown with the following displacements: 4.5m east, 7.5m 30° N of W, and 3.0m at 10° S of W. What is the resultant displacement?
1
Expert's answer
2020-10-21T09:53:55-0400

Using the following notions,



the vector of the corresponding displacements have the following coordinates:


"\mathbf{d}_1 = 4.5\times (1,0) = (4.5,0)\\
\mathbf{d}_2 =7.5\times (-\cos30\degree,\sin30\degree) \approx (-6.50,3.75)\\
\mathbf{d}_3 =3\times (-\cos10\degree,-\sin10\degree) \approx (-2.95,-0.52)\\"

The resultant displacement will be:


"\mathbf{d}_1 + \mathbf{d}_2+ \mathbf{d}_3 = \\
=(4.5 - 6.5 - 2.95, 0+3.75-0.52) = (-4.95,3.23) \space (m)"

Answer. "(-4.95,3.23) \space (m)".